OpenStudy (anonymous):

If Ф(x) = (a^x+a^-x) / 2 and ϴ(x) = (a^x-a^-x)/2 Prove that a)Ф(x+y)=Ф(x).Ф(y) + ϴ(x).ϴ(y) and b) ϴ(x+y)=Ф(x).ϴ(y) + Ф(y).ϴ(x)

OpenStudy (freckles):

a) for the first one I would play with the right hand side to show the left hand side

OpenStudy (freckles):

Same thing for b

OpenStudy (freckles):

Have you tried plugging in and doing the operations to see if you get the right hand side?

OpenStudy (aum):

\[ \phi(x) = \frac 12(a^x + a^{-x}) \\ \phi(x+y) = \frac 12 (a^{x+y} + a^{-{(x+y)} }) = \frac 12(a^x*a^y + a^{-x}*a^{-y}) \]Compute the right hand side separately and show they are equal.

OpenStudy (freckles):

\[\frac{a^x+a^{-x}}{2} \cdot \frac{a^y+a^{-y}}{2}+\frac{a^x-a^{-x}}{2}\cdot \frac{a^y-a^{-y}}{2}\]

OpenStudy (freckles):

try to multiply and add like terms if they should result

OpenStudy (freckles):

\[\frac{a^x+a^{-x}}{2} \cdot \frac{a^y+a^{-y}}{2}+\frac{a^x-a^{-x}}{2}\cdot \frac{a^y-a^{-y}}{2} \\ \frac{a^xa^y+a^xa^{-y}+a^{-x}a^y+a^{-x}a^{-y}}{4}+\frac{a^xa^y-a^xa^{-y}-a^{-x}a^{y}+a^{-x}a^{-y}}{4} \\\] combine like terms on top

OpenStudy (freckles):

from both fractions you should see a lot of cancellation

OpenStudy (freckles):

And you will see double of two terms

OpenStudy (freckles):

but guess what 2/4=1/2

OpenStudy (anonymous):

I think I lost myself =(

OpenStudy (freckles):

\[\frac{a^x+a^{-x}}{2} \cdot \frac{a^y+a^{-y}}{2}+\frac{a^x-a^{-x}}{2}\cdot \frac{a^y-a^{-y}}{2} \\ \frac{a^xa^y+a^xa^{-y}+a^{-x}a^y+a^{-x}a^{-y}}{4}+\frac{a^xa^y-a^xa^{-y}-a^{-x}a^{y}+a^{-x}a^{-y}}{4} \\ \frac{a^xa^y+\cancel{a^xa^{-y}}+\cancel{a^{-x}a^y}+a^{-x}a^{-y}}{4}+\frac{a^xa^y-\cancel{a^xa^{-y}}-\cancel{a^{-x}a^{y}}+a^{-x}a^{-y}}{4} \\ \]

OpenStudy (freckles):

you have a^xa^y+a^xa^y which is 2a^xa^y right?

OpenStudy (freckles):

and a^(-x)a^(-y)+a^(-x)a^(-y)=2a^(-x)a^(-y)

OpenStudy (freckles):

\[\frac{a^x+a^{-x}}{2} \cdot \frac{a^y+a^{-y}}{2}+\frac{a^x-a^{-x}}{2}\cdot \frac{a^y-a^{-y}}{2} \\ \frac{a^xa^y+a^xa^{-y}+a^{-x}a^y+a^{-x}a^{-y}}{4}+\frac{a^xa^y-a^xa^{-y}-a^{-x}a^{y}+a^{-x}a^{-y}}{4} \\ \frac{a^xa^y+\cancel{a^xa^{-y}}+\cancel{a^{-x}a^y}+a^{-x}a^{-y}}{4}+\frac{a^xa^y-\cancel{a^xa^{-y}}-\cancel{a^{-x}a^{y}}+a^{-x}a^{-y}}{4} \\ \frac{2a^xa^y+2a^{-x}a^{-y}}{4} \\ =2 \frac{a^{x+y}+a^{-x-y}}{4}\]

OpenStudy (freckles):

reduce the 2/4

OpenStudy (anonymous):

I was confused because your last answer already contain a step that would be enough to say that Ф (x + y) = Ф (x) .Ф (y) + Θ (x) .Θ (y) if the step it would be: (a^x+a^-x) /2 . (a^y + a^-y) / 2 + (a^x-a^-x)/2 + (a^y - a^-y) /2

OpenStudy (freckles):

i'm not sure what step you are talking about

OpenStudy (anonymous):

first step in answer above of "reduce the 2/4"